January 8, 2011

OBSERVERS: Jill Bechtold, David Sand, Marco Pedani, Michael Rutkowski, Russell Ryan

OPERATOR: David Gonzales Huerta

Our worries seem insignificant given the events of today.

Scattered clouds during the afternoon, but clear enough to try sky flats with Lucifer.

5pm. Opening. Very clear.

5:15pm. Starting sky flats. Br gamma in, waiting for sky to dim enough for K

5:35pm. Start K flats 4 sec exposures x 10

5:44pm FeII in, waiting for counts low enough to do H

5:48pm H band 2 sec x 10

5:52pm J band 10sec x 10

6pm. Going to point and collimate near c01_320_2512.txt for Fuyan. Waiting for sky to be dark enough to see pointing star.

6:15pm. Execute c01_320_2512 Waiting for it to be dark enough to see guide star well.

6:20pm Starting 30 minute integration. Ks Checked field with finder, we're good.

Seeing 1.2" Primary is 4 degrees warmer than ambient.

seeing 0.8 at end of exposure

7:04pm off to point up at MS0451

7:08pm Executing ms0451.Hacq.txt. Waiting for collimation on.

7:12 pm. Executing ms0451.H.txt. Waiting for collimation off. seeing is 2". elevation 37 degrees.

7:19pm. 3rd exposure: Images not round. Guider is oscillating, WFS is diverging.

7:22pm AWFUL. Stopped script. Will try to fix the collimation now. Removed all the corrections and will retune.

7:25pm re-executed ms0451.Hacq.txt.

wait for collimation is off.

7:32pm. Something is really messed up with collimation. Going to bright star.

7:42pm. Finally got ok collimation. Execute ms0451.Hacq.txt again. Wait for collimation is on.

elevation = 41 degrees.

7:47pm Execute ms0451.H.txt. Collimation still screwed up. Calling Doug Miller.

Aborting script.

7:57. Found a new guide star and PA=90 degrees. Will try this.

8:08pm Still collimation not quite converged. Exposure waiting for collimation.

Aborted and restarted, cleared collimation limits.

8:13 pm. Restarted ms0451.Hacq.txt, waiting for collimation.

8:16pm. Started ms0451.H.txt. seeing 1.1" star is round.

BOMBED after one exposure, couldn't offset.

8:19pm. Restarted ms0451.H.txt. 1.2" round images

LOST ONE HOUR FOR THIS CRAP.

TWO jitters put the GS off the field.

seeing .7"

8:47pm. Image FWHM .6". This is more like it.

9:02pm. Seeing .5" now

9:04pm. Executing ms0451.Jacq.txt, wait collimation off

9:06pm. Executing ms0451.J.txt, wait collimation off. Seeing .5"

9:52pm. Executing sdss0405.Hacq.txt with wait collimation on. Found guide star.

9:55pm. Executing sdss0405.H.txt with wait coll on. Seeing .7", humidity rising, 63%. Checked field with K-band image, looks good.

10:55pm. Script stopped after 40 minutes of exposure time, lost guide star, clouds. Sky looks bad.

11:01pm. Closing the dome.

11:40pm open again, off to try sdss0854. Pointing first.

11:42pm. Execute sdss0854.Hacq.txt

11:45pm. Execute sdss0854.H.txt. Elevation 45 degrees. Cloudy. Took 4 minutes for collimation to settle down. seeing .8"

12:27pm. Seeing deteriorating, 1.2"

1:23am. Execute restof_0854.H.txt. Still murky out. Seeing .8" in H.

1:34am. Seeing .7" Clouds getting thicker.

1:44am Clouds. Script stopped after 14 dithers.

1:45am. reexecuted restof_0854.H.txt. Clouds coming and going. Seeing .9"

3:19am. Off to point up at MACS1149

3:22am. Execute macs1149.Kacq.txt, wait collim. on

3:26am Start macs1149.K.txt. Seeing .9"

Stopped with 3 minutes left, rotators in fault. can't preset telescope

4:22am Fixed after call to software guy. Rebooted pcs. macs1149.Hacq.txt

no good. won't preset

4:26am. Re-try, preset successful

4:29am. macs1149H.txt

1.2" seeing

Stop after 27 dithers -- dewpoint alarm went off.

5:09am. Closing dome.

high humidity the rest of the night.

Took darks.

Topic revision: r4 - 09 Jan 2011, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback